40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Feb 25 15:30:29 2013, yuta, Update, Alignment, Y arm locked, both colors 
    Reply  Mon Feb 25 17:58:28 2013, rana, Update, Alignment, Y arm locked, both colors 
       Reply  Mon Feb 25 22:42:32 2013, yuta, Summary, Alignment, current IFO situation IPANGdrift.png
    Reply  Mon Feb 25 21:25:14 2013, yuta, Update, Alignment, both arms locked in IR QUAD1_1045890588.png
       Reply  Tue Feb 26 01:49:27 2013, Jenne, Update, Locking, PRMI locked 
          Reply  Tue Feb 26 02:04:51 2013, Koji, Update, Locking, PRMI locked 
             Reply  Tue Feb 26 11:55:11 2013, Manasa, Update, Locking, FPMI locked 
                Reply  Tue Feb 26 15:32:47 2013, Jenne, Update, Locking, FPMI locked 
             Reply  Wed Feb 27 01:09:53 2013, yuta, Update, Locking, PRMI sideband/carrier locked 6x
                Reply  Wed Feb 27 02:44:39 2013, Jenne, Update, Locking, DRFPMI flashes 
                   Reply  Wed Feb 27 20:10:43 2013, rana, Update, Locking, DRFPMI flashes 
Message ID: 8166     Entry time: Tue Feb 26 02:04:51 2013     In reply to: 8165     Reply to this: 8170   8177
Author: Koji 
Type: Update 
Category: Locking 
Subject: PRMI locked 

One of the biggest issues we had was that any Q signals (i.e. the quadrature where PRCL is absent.) of REFL11/33/55/(165) haven't been consistent each other.
i.e. We never had reliable lock of MICH with REFL_any_Q, regardless of the resonant condition. This is definitely one of the things to be tried in order to prepare for the full lock.

We don't trust any demodulation phases of any PDs any more as the previous PRC mode (or say, absence of the stable mode) was unreasonable to determine any of the demodulation phases.

I remember that the POP DC saturates at 27000. You may need to reduce the gain switch again.

The AS OSA and/or POP BBPD would be useful for the sideband PR gain estimation.

ELOG V3.1.3-